# Describe-and-Dissect: Interpreting Neurons in Vision Networks with Language Models

### v1

@@ A.4.2 Ablation: Image Captioning with Fixed Concept Set
To analyze the importance of using a generative image-to-text model, we explore instead utilizing fixed concept sets with CLIP (Radford et al., 2021) to generate descriptions for each image instead of BLIP, while the rest of the pipeline is kept the same (i.e. using GPT to summarize etc). For the experiment, we use CLIP-ViT-B/16, where we define $L(\cdot)$ and $E(\cdot)$ as text and image encoders respectively. From the initial concept set $\mathcal{S}=\{t_{1},t_{2},...\}$, the best concept for image $I_{m}$ is defined as $t_{l}$, where $l=\text{argmax}_{i}(L(t_{i})\cdot E(I_{m})^{\top})$. Following CLIP-dissect (Oikarinen & Weng, 2023), we use $\mathcal{S}$ = 20k [^2] Source: https://github.com/first20hours/google-10000-english/blob/master/20k.txt($20,000$ most common English words) and $\mathcal{D}_{probe}=\text{ImageNet}\cup\text{Broden}$.
To compare the performance, following (Oikarinen & Weng, 2023), we use our model to describe the final layer neurons of ResNet-50 (where we know their ground truth role) and compare descriptions similarity to the class name that neuron is detecting, as discussed in Section 4.2.1. Results in Table 8 show that both methods perform similarly on the FC layer. In intermediate layers, we notice that single word concept captions from 20k significantly limit the expressiveness of DnD, and that having generative image descriptions is important for our overall performance. Qualitative examples are shown in Figure 16 and Figure 17 showcases some failures of using CLIP descriptions.
@@
Table 8: Mean FC Layer Similarity of CLIP Captioning. Utilizing a fixed concept set ($\mathcal{S}$ = 20k) to caption activating images via CLIP (Radford et al., 2021), we compute the mean cosine similarity across fully connected layers of RN50. Final labels that do not express concepts (see Figure 17) are given a similarity score of 0. With CLIP cosine similarity, mpnet cosine similarity, and BERTScore, we find the performance of DnD w/ CLIP Captioning is slightly worse than BLIP generative caption.
Metric / Methods Describe-and-Dissect (Ours) DnD w/ CLIP Captioning % Decline
CLIP cos 0.7598 0.7583 0.197%
mpnet cos 0.4588 0.4465 2.681%
BERTScore 0.8286 0.8262 0.290%
@@ A.4.3 Ablation: Image-to-Text Model
In light of advancements in image-to-text models, we compare BLIP (Li et al., 2022) to a more recently released model, BLIP-2 (Li et al., 2023). Unlike BLIP, BLIP-2 leverages frozen image encoders and LLMs while introducing Querying Transformer to bridge the modality gap between two models. We experiment with using BLIP-2 as the image-to-text model and quantitatively compare with BLIP by computing the mean cosine similarity between the best concept chosen from Concept Selection. As discussed in the Appendix A.3, CLIP-ViT-B/16 cosine similarity is a stronger indicator of conceptual connections than all-mpnet-base-v2 similarity for generative labels. Accordingly, CLIP-ViT-B/16 cosine similarity is used as the comparison metric. We evaluate on 50 randomly chosen neurons from each intermediate layer of RN50 and results of the experiment are detailed in Table 9. From Table 9, BLIP and BLIP-2 produce highly correlated concepts across all four layers of RN50 and a 87.0% similarity across the entire network.
@@
We also show examples of similar description produced in Figure 18. Somewhat counter-intuitively, our qualitative analysis of neurons in preliminary layers of RN50 reveals examples where BLIP-2 fails to detect low level concepts that BLIP is able to capture. Such failure cases limit DnD accuracy by generating inaccurate image captions and adversely affecting the ability of GPT summarization to identify conceptual similarities between captions. We show two such examples in below Figure 19. In general this experiment shows that our pipeline is not very sensitive to the choice of image-captioning model.
@@
Figure 18: Examples of Similar BLIP and BLIP-2 Labels. BLIP and BLIP-2 detect similar concepts across all layers of RN50. Two such examples are detailed above.
@@
Figure 19: Examples of BLIP-2 failure cases. BLIP-2 overlooks crucial low level concepts in early layers of RN50. Potential Candidate Concepts generated are vague and spuriously correlated, yielding poor results in the final DnD label.
A.4.4 Ablation: Effects of GPT Concept Summarization
@@
  • Concept specific to a single image. Without the summarization step, concepts only describe a single image and fail to capture the shared concept across images. We can see for example that the neuron in Figure 20b gets only described as a ”pile of oranges”, which only applies to the fourth image and misses the overall theme of orange color.
@@
Figure 20: Failure Cases in GPT-Ablation Study. Figure a illustrates potential failure cases of DnD without GPT summarization. BLIP produces illogical image captions that hinder overall interpretability. Figure b shows an additional failure case where the omission of concept summarization causes individual image captions to generalize to the entire neuron.
A.4.5 Effects of Concept Selection
Method / Layer Layer 1 Layer 2 Layer 3 Layer 4 All Layers
DnD (w/o Best Concept Selection) 3.54 3.77 4.00 4.02 3.84
DnD (full pipeline) 3.54 4.00 4.24 4.13 3.97
Using the same setup as section A.1.4, we conducted an ablation study to determine the effectiveness of our Best Concept Selection (step 3) on the pipeline. Table 10 shows the effect of Best Concept Selection on the overall accuracy of DnD. We can see DnD performance is already high without Best Concept Selection, but Concept Selection further improves the quality of selected labels Layers 2 through Layer 4, while having the same performance on Layer 1. One potential explanation is due to Layer 1 detecting more limited lower level concepts – there is less variance in candidate descriptions identified in Concept Generation (step 2), resulting in similar ratings across the set of candidate concepts $T$. We can see some individual examples of the improvement Concept Selection provides in Figure 21, with the new labels yielding more specific and accurate descriptions of the neuron. For example Layer 2 Neuron 312 becomes more specific colorful festive settings instead of generic Visual Elements. We include the same Figure 3 from Section 4.5 for reader convenience.
((a))
((b))
Figure 21: Concept Selection (Step 3) supplements Concept Generation (Step 2) accuracy. We show that concept selection improves Concept Generation by removing vague and spuriously correlated concepts. Figure 21 is moved from Section 4.5 for reader convenience.
A.5 Multiple Labels
@@
As discussed in the Appendix A.6, many neurons in intermediate layers can be considered ”polysemantic”, meaning that they encode multiple unrelated concepts. Our primary pipeline only produces one label, and though this label can encapsulate more than one concept, providing multiple labels can better account for issues of polysemanticity. We accomplish this by taking the top candidate concepts selected by Concept Selection as our final labels. If the final labels have a CLIP cosine similarity exceeding 0.81, we only take the top labels and eliminate the others from the set of final labels. This allows use to bring out accurate labels that account for distinct concepts prevalent in the neuron. We can see this in Figure 22. For Neuron 508 from Layer 2, DnD captures not only the polka dot and metal texture concepts of the neuron, but adds that the textures are primarily black and white. For Neuron 511 from Layer 3, DnD labels the neuron as detecting not only interior elements, but also specifying that these elements are mostly black and white.
@@
A.6 Limitations
@@
One limitation of Describe-and-Dissect is the relatively high computational cost, taking on average about 38.8 seconds per neuron with a Tesla V100 GPU. However, this problem can be likely well-addressed as the current pipeline has not been optimized for speed-purposes yet (e.g. leveraging multiple GPU etc). Another potential limitation of our method is that since it first converts the images into text and then generates its labels based on the text, DnD likely cannot detect purely visual elements occasionally found in lower layers of networks, such as contrast or spatial elements. Additionally, our method only takes into account the top $k$ most highly activating images when generating labels. Excluding the information found in lower activating images may not allow our method to gain a full picture of a neuron. However, this problem of only focusing on top activations is shared by all existing methods and we are not aware of good solutions to it. Finally, DnD is limited by how well the image-to-text, natural language processing, and text-to-image models are able to perform. However, this also means that future innovations in these types of Machine Learning models can increase the performance of our method.
@@
Polysemantic neurons: Existing works (Olah et al., 2020; Mu & Andreas, 2020; Scherlis et al., 2023) have shown that many neurons in common neural networks are polysemantic, i.e. represent several unrelated concepts or no clear concept at all. This is a challenge when attempting to provide simple text descriptions to individual neurons, and is a limitation of our approach, but can be somewhat addressed via methods such as adjusting DnD to provide multiple descriptions per neuron as we have done in the Appendix A.5. We also note that due to the generative nature of DnD, even its single labels can often encapsulate multiple concepts by using coordinating conjunctions and lengthier descriptions. However, polysemantic neurons still remain a problem to us and other existing methods such as (Bau et al., 2017) (Hernandez et al., 2022) (Oikarinen & Weng, 2023). One promising recent direction to alleviate polysemanticity is via sparse autoencoders as explored by (Bricken et al., 2023).
@@
Challenges in comparing different methods: Providing a fair comparison between descriptions generated by different methods is a very challenging task. The main method we (and previous work) utilized was displaying highly activating images and asking humans whether the description matches these images. This evaluation itself has some flaws, such as only focusing on the top-k activations and ignoring other neuron behavior. In addition, the question of what are the highly activating images and how to display them is a surprisingly complex one and requires multiple judgement calls, with different papers using different methods with little to no justification for their choices. First, which images are most highly activating. When our neuron is a CNN channel, its activation is 2D – sorting requires summarizing to a scalar, typically using either max or average pooling. In our experiments we used average, but max gives different sets of results. Second, the choice of probing dataset is important, as it will affect which images are displayed. Finally choosing if/how to highlight the highly activating region to raters will have an effect. We chose not to highlight as we crop the highly activating regions which can provide similar information. If highlighting is used, choices like what activation threshold to use for highlights and how to display non-activating regions will also affect results. Different choices for all these parameters will generate varying results. Consequently, it is hard to provide fully fair comparisons, so results such as Section 4.3 should be taken with a grain of salt.

### v2

@@ A.4.2 Ablation: Image Captioning with Fixed Concept Set
In Section 4.4.1, we explored using fixed concept sets with CLIP (Radford et al., 2021) to generate descriptions for highly activating images rather than BLIP. We present qualitative examples of fixed concept captioning below in Figure 20. Additionally, we also note particular failure cases caused by a lack of expressiveness encapsulated in static concepts, as demonstrated in Figure 21.
@@ A.4.3 Ablation: Image-to-Text Model
In light of advancements in image-to-text models, we compare BLIP (Li et al., 2022) to a more recently released model, BLIP-2 (Li et al., 2023). Unlike BLIP, BLIP-2 leverages frozen image encoders and LLMs while introducing Querying Transformer to bridge the modality gap between two models. We experiment with using BLIP-2 as the image-to-text model and quantitatively compare with BLIP by computing the mean cosine similarity between the best concept chosen from Concept Selection. As discussed in the Appendix A.3, CLIP-ViT-B/16 cosine similarity is a stronger indicator of conceptual connections than all-mpnet-base-v2 similarity for generative labels. Accordingly, CLIP-ViT-B/16 cosine similarity is used as the comparison metric. We evaluate on 50 randomly chosen neurons from each intermediate layer of ResNet-50 and results of the experiment are detailed in Table 10. From Table 10, BLIP and BLIP-2 produce highly related concepts across all four layers of ResNet-50 and a 87.0% similarity across the entire network.
@@
We also show examples of similar description produced in Figure 22. Somewhat counter-intuitively, our qualitative analysis of neurons in preliminary layers of ResNet-50 reveals examples where BLIP-2 fails to detect low level concepts that BLIP is able to capture. Such failure cases limit DnD accuracy by generating inaccurate image captions and adversely affecting the ability of GPT summarization to identify conceptual similarities between captions. We show two such examples in below Figure 23. In general this experiment shows that our pipeline is not very sensitive to the choice of image-captioning model.
@@
Figure 22: Examples of Similar BLIP and BLIP-2 Labels. BLIP and BLIP-2 detect similar concepts across all layers of ResNet-50. Two such examples are detailed above.
@@
Figure 23: Examples of BLIP-2 failure cases. BLIP-2 overlooks crucial low level concepts in early layers of ResNet-50. Potential Candidate Concepts generated are vague and spuriously correlated, yielding poor results in the final DnD label.
A.4.4 Ablation: Multimodal GPT
To justify using two large models instead of one during Step 2 in our pipeline, we conduct an additional experiment by exchanging BLIP + GPT in Step 2 of DnD with the image caption capabilities of GPT-4o. Results in Figure 24 show that due to the large overlap of the GPT models, performance is similar in many examples. However, the original DnD pipeline seems to be better than DnD w/ GPT-4o when BLIP’s image captioning capabilities outperform that of GPT-4o. Though including two foundation models in Step 2 may reduce the flexibility of our pipeline, it can increase performance by specifying fine-tuned models for each designated task.
Figure 24: Neuron examples of GPT-4o replacing BLIP and GPT-3.5-Turbo in the DnD pipeline.
A.4.5 Ablation: Effects of Large-Language-Model Choice
To test the robustness of our pipeline at the concept summarization stage, we swap out GPT-3.5 Turbo for GPT-4 Turbo and LLaMA2 (Touvron et al., 2023). For reference, our paper utilizes GPT-3.5 Turbo for all of its primary experiments rather than GPT-4 Turbo because GPT-4 Turbo has 20x the cost of GPT-3.5 Turbo. At the time of developing our pipeline, GPT-3.5 Turbo was the most advanced GPT model provided by OpenAI. LLaMA2, although cost free, is restricted to the public and requires specialized access for use in research.
Our experiment follows the precedence of the setups of our previous experiments in our paper and of the setups utilized in prior works. We evaluate 26 neurons on the FC Layer of ResNet-50 and compare to the ground truth labels provided by the classes. In this experiment, we can see that utilizing other models actually significantly increases DnD’s performance in all metrics. As shown in Table 11, not only is DnD able to keep up its performance with other models but also improve with new and more advanced ones.
Table 11: Comparison of LLMs on label quality. Our experiment shows that GPT-4 Turbo and LLaMA2 significantly increase the cosine similarity score between DnD labels and ground truth classes on FC layer neurons, suggesting the ability for our framework to incorporate new advancements in LLMs.
Metric DnD DnD w/ GPT-4 Turbo DnD w/ LLaMA2
CLIP cos 0.6377 0.7607 0.7407
mpnet cos 0.1403 0.4506 0.4726
BERT Score 0.8172 0.8340 0.8105
A.4.6 Ablation: Effects of GPT Concept Summarization
@@
  • Concept specific to a single image. Without the summarization step, concepts only describe a single image and fail to capture the shared concept across images. We can see for example that the neuron in Figure 25b is described as a "pile of oranges", which only applies to the fourth image and misses the overall theme of orange color.
@@
Figure 25: Failure Cases in GPT-Ablation Study. Figure (a) illustrates potential failure cases of DnD without GPT summarization. BLIP produces illogical image captions that hinder overall interpretability. Figure (b) shows an additional failure case where the omission of concept summarization causes individual image captions to generalize to the entire neuron.
A.5 Use Case: Fixed Classifier
The Tile2Vec model evaluated contains two parts: 1. A ResNet model to compute embeddings of land tiles, 2. A random forest classifier to predict the land coverage of the output embedding. We use the originally proposed Tile2Vec pipeline by first pruning neurons in the ResNet model and then training the classifier on embeddings of the pruned model. Because we retrain the classifier on the pruned pipeline, the difference in accuracy using DnD pruning and that of random pruning is fairly minor. In this section, we instead evaluate the Tile2Vec model using a fixed classifier. We train the classifier using embeddings from the full unpruned pipeline and evaluate by fixing the classifier while pruning the ResNet model only. We observed a moderate (13.63%) increase in speed after performing pruning on the first layer of Tile2Vec ResNet-18 and further gains are possible if running inference more optimized for sparse models. Table 12 shows results using a similar percentage of pruned neurons as Table 6.
Table 12: Pruning uninterpretable neurons in Tile2Vec ResNet18. Due to GPT’s generative captioning, the similarity between neuron labels can vary slightly between runs, causing slightly different number of neurons pruned.
Layer % of Neurons Pruned Random Pruning Acc. (%) Avg. Acc. (%)
No pruning 0.00 71.63
All pruned 100.00 35.96
Layer 1 23.44 45.0 51.0
Layer 2 50.0 47.0 49.5
Layer 3 26.95 66.6 70.5
Layer 4 50.78 49.5 56.0
Layer 5 56.64 56.0 60.5
A.6 Additional Use Case
To showcase a potential use case for neuron descriptions (and provide another way to quantitatively compare explanation methods), we experimented with using neuron descriptions to find a good classifier for a class missing from the training set. Our setup was as follows: we explained all neurons in Layer 4 of ResNet-50 (ImageNet) using different methods. We then wanted to find neurons in this layer that could serve as the best classifiers for an unseen class, specifically the classes in CIFAR-10, CIFAR-100, and Places365 datasets. Although there is some overlap, these classes are typically much more broad than ImageNet classes. To find a neuron to serve as a classifier, we locate the neuron whose description is closest to the CIFAR/Places365 class name in a text embedding space. We then measure how well that neuron (its average activation) performs as a single class classifier on the respective validation dataset, measured by area under ROC curve. For cases where multiple neurons share the closest description, we average the performance of all neurons with that description.
Results are shown in Table 13. We can see DnD performs quite well, reaching AUROC values around 0.74, while MILAN performs much worse. We believe this may be because MILAN descriptions are very generic (likely caused by noisy dataset as discussed in Section 4.3.1), making it hard to find a classifier for a specific class. We believe this is a good measure of explanation quality, as different methods are dissecting the same network, and even if no neurons exist that can directly detect a class, a better method should find a closer approximation.
MILAN DnD (Ours)
CIFAR10 0.5906 0.7036
CIFAR100 0.6514 0.7396
Places365 0.650 0.709
A.7 Multiple Labels
@@
As discussed in the Appendix A.10, many neurons in intermediate layers can be considered "polysemantic", meaning that they encode multiple unrelated concepts. Our primary pipeline only produces one label, and though this label can encapsulate more than one concept, providing multiple labels can better account for issues of polysemanticity. We accomplish this by taking the top candidate concepts selected by Concept Selection as our final labels. If the final labels have a CLIP cosine similarity exceeding 0.81, we only take the top labels and eliminate the others from the set of final labels. This allows use to bring out accurate labels that account for distinct concepts prevalent in the neuron. We can see this in Figure 26. For Neuron 508 from Layer 2, DnD captures not only the polka dot and metal texture concepts of the neuron, but adds that the textures are primarily black and white. For Neuron 511 from Layer 3, DnD labels the neuron as detecting not only interior elements, but also specifying that these elements are mostly black and white.
@@
A.8 Diversity Analysis
To more deeply understand the effectiveness and outcomes of our method, we perform an experiment analyzing the relationship between similar highly activating images and similar labels. We follow the setup in Section 4.2. Over the entirety of the FC Layer of ResNet-50 (1000 neurons), we took random pairs of neurons from those 1000, making 500 pairs in total. We then calculated the similarity between each pair’s top 10 activating images and calculated the CLIP cosine similarity between each pair’s labels outputted by DnD. Finally, we calculated the correlation coefficient between image similarity and label similarity with each pair being a datapoint and found that value to be 0.4406. We can see there is a positive correlation between similar highly activating images and similar DnD labels. In general, DnD provides similar labels for neurons with similar highly activating images while more varying labels for neurons with differing highly activating images.
A.9 Failure Cases
In this section we analyze a few examples in Figure 27 that can cause DnD to produce suboptimal neuron labels. Neuron 106 from Layer 3 exemplifies how DnD can occasionally focus on one specific element from the highly activating image set when the true concept is less obvious. Neuron 193 from Layer 3 shows how when given an uninterpretable or polysemantic neuron, DnD defaults to a generic label like "patterned items." These neurons are further discussed in sections A.7 and A.10. Neuron 515 from Layer 3 demonstrates how DnD may produce a more generic and altered version of the true concept when the true concept is more specific. We can see that the true concept represented by the neuron is likely something along the lines of "animal ears," but DnD is not able to identify the ears and adds in an incorrect portion of "resting" to the label. Finally, neuron 1880 from Layer 4 showcases how it’s hard for DnD to detect more spacial and abstract representations of the concept. We can see that the true concept should be likely along the lines of "line patterns" but since the lines are embedded into other concepts like nature elements, DnD isn’t able to detect it.
Figure 27: Failure cases of DnD. We showcase specific neuron examples from dissecting ResNet-50 that exemplify instances where DnD can produce ambiguous or inaccurate labels.
A.10 Limitations
@@
One limitation of Describe-and-Dissect is the relatively high computational cost, taking on average about 38.8 seconds per neuron with a Tesla V100 GPU. However, this problem can be likely well-addressed as the current pipeline has not been optimized for speed-purposes yet (e.g. leveraging multiple GPU etc). Another potential limitation of our method is that since it first converts the images into text and then generates its labels based on the text, DnD likely cannot detect purely visual elements occasionally found in lower layers of networks, such as contrast or spatial elements. Additionally, our method only takes into account the top $k$ most highly activating images when generating labels. Excluding the information found in lower activating images may not allow our method to gain a full picture of a neuron. However, this problem of only focusing on top activations is shared by all existing methods and we are not aware of good solutions to it. Finally, DnD is limited by how well the image-to-text, natural language processing, and text-to-image models are able to perform. On the other hand, this also means that future innovations in these types of Machine Learning models can increase the performance of our method.
@@
Polysemantic neurons: Existing works Olah et al. (2020); Mu & Andreas (2020); Scherlis et al. (2023) have shown that many neurons in common neural networks are polysemantic, i.e. represent several unrelated concepts or no clear concept at all. This is a challenge when attempting to provide simple text descriptions to individual neurons, and is a limitation of our approach, but can be somewhat addressed via methods such as adjusting DnD to provide multiple descriptions per neuron as we have done in the Appendix A.7. We also note that due to the generative nature of DnD, even its single labels can often encapsulate multiple concepts by using coordinating conjunctions and lengthier descriptions. However, polysemantic neurons still remain a problem to us and other existing methods such as Bau et al. (2017), Hernandez et al. (2022), and Oikarinen & Weng (2023). One promising recent direction to alleviate polysemanticity is via sparse autoencoders as explored by Bricken et al. (2023).
Out-of-Distribution Synthetic Images: While generated synthetic images in Step 3 (Best Concept Selection) may be out of distribution, we note that an explanation should explain the behavior of the model on all inputs, not just in-distribution examples. A neuron explained by “car” should activate on all kinds of cars, not just in-distribution ones, otherwise, a better explanation would be more specific.
@@
Challenges in comparing different methods: Providing a fair comparison between descriptions generated by different methods is a very challenging task. The main method we (and previous work) utilized was displaying highly activating images and asking humans whether the description matches these images. This evaluation itself has some flaws, such as only focusing on the top-k activations and ignoring other neuron behavior. In addition, the question of what are the highly activating images and how to display them is a surprisingly complex one and requires multiple judgment calls, with different papers using different methods with little to no justification for their choices. First, which images are most highly activating. When our neuron is a CNN channel, its activation is 2D – sorting requires summarizing to a scalar, typically using either max or average pooling. In our experiments we used average, but max gives different sets of results. Second, the choice of probing dataset is important, as it will affect which images are displayed. Finally choosing if/how to highlight the highly activating region to raters will have an effect. We chose not to highlight as we crop the highly activating regions which can provide similar information. If highlighting is used, choices like what activation threshold to use for highlights and how to display non-activating regions will also affect results. Different choices for all these parameters will generate varying results. Consequently, it is hard to provide fully fair comparisons, so results such as Section 4.3 should be taken with a grain of salt.
A.11 Broader Impact
Our work aims to improve our understanding of machine learning models, which we expect to have positive societal impact as more understanding will decrease the chance of harm from deploying unsafe or unreliable models. However, one potential downside may be that if explanations look convincing but are not reliable enough, they may cause a user to place unwarranted trust on an unreliable model. Our methods can be used to help reveal and mitigate biases in existing vision models by, for example, identifying neurons detecting sensitive attributes such as skin-color.